Calculate the density (at STP) of:
(i). hydrogen, H2

(ii). nitrogen, N2

(iii). butane, C4H10
1
Expert's answer
2020-09-23T04:46:17-0400

Assuming that we have 1 mole of each gas at STP, we will use Molar mass and Molar volume to calculate density of each gas.

Molar Volume of any gas at STP is 22400mL

(i) Hydrogen, H2

Molar mass=2.016g/mol

Molar volume=22400mL/mol

"Density=\dfrac{mass}{volume}=\dfrac{2.016g/mol}{22400mL/mol}"


= 9.0x10-5g/mL


(ii) Nitrogen, N2

Molar mass=28.0134g/mol

molar volume=22400mL

"Density=\dfrac{28.0134g/mol}{22400g/mol}=" 1.25x10-3g/mL


(iii) Butane, C4H10

Atomic mass=(C=12.0, H=1.008)


Molar mass= 4(12.0) + 10(1.008) = 58.08g/mol

Molar volume=22400mL

"Density=\dfrac{58.08g/mol}{22400mL/mol} =" 2.59x10-3g/mL